# SeBS-Flow: Benchmarking Serverless Cloud Function Workflows

### v2

@@ VII-B RQ1: Runtime Differences among Platforms
Figure 7: Runtime of benchmark applications on AWS Step Functions, GC Workflows, and Azure Durable, burst invocations.
VII-C RQ2: Causes for Runtime and Stability Differences.
According to our results, AWS and Google Cloud provide a performance-reliable workflow service, whereas the variability is considerably higher on Azure. Thus, we split the runtime into two components: the critical path $T_{C}$ computed as the sum of all states’ maximum runtime within one phase, and the overhead $T_{O}$ caused by the scheduling and data movement conducted by the cloud workflow service. We calculate the overhead by subtracting the critical path from the total runtime. Figure 8 presents the critical path and overhead for all benchmarks. Azure’s runtime is dominated by scheduling overhead: For example, the overhead of the ExCamera benchmark is, on average, 495.5s, more than 36$\times$ as long as its critical path of 13.5s. The ML benchmark incurs the least overhead of 5$\times$ the length of its critical path. Also, Azure’s critical path is very fast across all benchmarks, demonstrating the fastest critical path for ExCamera, MapReduce, and Machine Learning. Google Cloud, however, has the slowest critical path throughout the entire benchmark suite. In summary, orchestration overhead causes long runtimes and performance variances on Azure. For AWS and Google Cloud, however, the critical path varies.
@@
Figure 8: Critical path (opaque) and overhead (hatched) of different benchmarks on considered platforms, burst invocations.
VII-C1 RQ2.1 Sources of Overhead
@@
The data downloaded from the object storage differs between benchmarks (Table IV, p. IV), with hundreds of megabytes in ExCamera, 1000Genomes, and Video Analysis. These benchmarks experience the highest relative overhead of 36.7$\times$, 10$\times$, and 14.95$\times$ their critical paths on Azure. To verify that this correlation is indeed causation, we execute a microbenchmark evaluating the cloud storage I/O performance. We invoke 20 functions in parallel where each attempts to download a file of size $D$ from the storage. Figure 9(a) shows that the overhead remains stagnant for AWS at around one second and nearly stagnant on Google Cloud, increasing a bit for downloads larger than 1MB. On Azure, we observe an overhead of almost 149 and 4.9 seconds for 128 and 1 MB files, respectively. This can account for a significant part of the large overhead measured on Azure Durable.
@@
(b) Invocation latency,
$2^{5}\leq M<2^{18}$, 256MB, warm invocations.
Figure 9: Analysis of different sources of overhead.
@@
Another potential source of overhead are parallel invocations within a benchmark. Benchmarks with the highest degree of parallelism – ExCamera and 1000Genomes – show the largest overheads of Azure. We test this by executing a microbenchmark that spawns $N$ functions in parallel, each one sleeping for $T$ seconds, and start 30 such invocations concurrently. Figure 10 shows the relative overhead of the actual runtime compared to the function execution time. AWS functions demonstrate modest overhead, with largest values for the shortest duration. GC functions present a larger slowdown that increases with the number of parallel tasks. There, the system puts a cap on scaling up and reuses containers, as 30 invocations with $N=2,T=1$ start 60 different function containers on AWS, but only 30 on Google Cloud. On the other hand, Azure experiences an order of magnitude larger overhead that increases with the parallelism factor but does not seem to be correlated to the function runtime.
@@
Figure 10: The overhead of parallel sleep microbenchmark, $2\leq N\leq 16$, $1\leq T\leq 20$, 256MB, burst invocations.
@@
Figure 11: Scaling profiles: the number of distinct containers used for 30 consecutive workflow invocations.
@@ Conclusions
The microbenchmarks demonstrate that a significant part of the overhead observed on Azure originates from the parallel schedules and storage IO. Another potential source is dynamic orchestration. A statically scheduled system could optimize function placement, data prefetching, and scalability by using a priori knowledge.
VII-C2 RQ2.2 Critical Path Discrepancy
@@
The cloud provider controls the CPU allocation to a serverless function, either in relation to the memory configuration on AWS and GCP [52, 28], or in an undisclosed fashion on Azure. We use the selfish detour benchmark to quantify OS noise [53], which allows us to estimate how long the function is suspended by the OS, which in turn approximates the vCPU timeshare. The benchmark runs a tight loop and records the event that one iteration took significantly more cycles than expected $N$ times. The magnitude and frequency of these events characterize the suspension and noise. We deploy a workflow with a single function executing the benchmark, invoke it 30 times concurrently, collect $N=5000$ events, and sample warm invocations to obtain consistent results. Figure 13(a) compares the relative to the expected suspension time according to the cloud documentation. We observe less noise on Google Cloud when compared to AWS, with more than 20% difference on 1024MB memory. We normalize the critical path per platform using the following approximation: given a function with memory configuration $M$, we represent the relative duration of function suspension as $S_{M}$ and compute the normalized critical path $T^{\prime}_{C}=T_{C}*\left(1-S_{M}\right)$. We observe the largest relative discrepancy on two benchmarks, MapReduce (Figure 13(b)) and Machine Learning (Figure 13(c)). The overall trend observed in Section VII-B remains unchanged: Google Cloud demonstrates the longest critical path duration. The suspension time explains the shorter critical path on Azure when testing low-memory function configurations on AWS and GCP: Azure functions receive larger CPU allocations.
Cold Starts.
@@
Cold invocations add significant overhead to the function execution [11]. Table V shows the frequency of cold starts in our measurements, with cold starts identified using the containerID (see Section IV-C). Azure Durable performs significantly better, likely because function apps on Azure can hold many invocations concurrently [11]. While the low scalability causes high orchestration overheads, it benefits the computations by putting them in warm containers. Figure 12 shows the impact of cold starts on the critical path and overhead. Due to the high percentage of cold starts in our measurement data, we collected another 60 workflow invocations for AWS and GCP with at least one warm function and show the critical path for the resulting completely warm invocations. Google Cloud and AWS functions perform up to 2.0×2.0$\times$2.0 × and 4.5×4.5$\times$4.5 × better, respectively, achieving almost the same performance as Azure. Thus, cold starts are a major factor influencing the slowdown and performance instability observed in many benchmarks.
@@
TABLE V: Relative #cold starts and #state transitions.
@@
Figure 12: Critical path (opaque) and overhead (hatched) of warm and cold invocations.
@@
Figure 13: Analysis of OS noise.
@@
Figure 14: Scalability of the 1000Genome workflow.
VII-D RQ3: Usability for Scientific Workflows
@@
There is rising interest in the scientific community to use serverless solutions [41], accompanied by experimentation with serverless offerings of the platforms [54] and management systems for serverless execution of workflows [55, 56, 57, 58]. However, they do not consider the workflow orchestration systems the cloud platforms offer. We use the scientific benchmark 1000Genome to compare cloud services and the HPC system Ault using nodes equipped with Intel(R) 6154@3.00GHz CPU, repeating measurements five times.
@@
First, we compare the runtime of the total workflow, as shown in Figure 13(d). While the workflow execution time is, on average, 457.7s and 259.8s on GCP and AWS, respectively, the execution takes only 7.7s on Ault. GCP exhibits a coefficient of variation of 12.2%, while AWS has a coefficient of variation of only 3.3% - even lower than 4.1% on Ault. Interestingly, I/O takes less than one second on AWS, meaning that the computation is slower in the cloud. Then, we compare the scaling behavior of the different platforms for the individuals task of the workflow. We employ strong scaling, i.e., adding more jobs while keeping the size of the input file the same, resulting in smaller chunks per job. Figure 13(e) shows the speedup of 1.96 and 1.95 on AWS, 1.91 and 1.95 on GCP, and 1.51 and 1.24 on Ault for 10 and 20 jobs, respectively. The cloud platforms achieve a nearly-optimal speedup, which is not surprising given the high overhead for the baseline execution.
VII-E RQ4: Pricing
@@
We compare the average cost of executing a workflow and estimate the prices, as shown in Table III, p. III. Functions invoked during the execution of a workflow are billed based on the integral of memory and duration. Figure 15 visualizes the cost of workflow execution split into two groups: function execution (opaque) and the cost of orchestrating the state machine (hatched). Note that, due to Azure’s billing and measurement system, we could only retrieve an average cost value over all workflow invocations. Even though the Trip Booking benchmark is a simple pipeline with error catching, running it with workflow orchestration still adds significant state transition costs. Azure is the most expensive service for the 1000Genome benchmark. Google Cloud is the most expensive for MapReduce due to the high number of state transitions. AWS Step Functions are the most expensive solution for the other four benchmarks because functions cost 6.7×6.7$\times$6.7 × more for computation than Google Cloud Functions. The price charged for state transitions is nearly identical between AWS and Google Cloud, even though AWS charges 2.5×2.5$\times$2.5 × more: the AWS state language requires fewer states to implement the benchmarks (Table V).
@@
In addition to execution and orchestration costs, workflows can generate charges when accessing the object and NoSQL storage. In all three clouds, the prices of read and write operations on the object storage are exactly the same. However, the billing models for key-value storage differ: DynamoDB charges for operations according to the amount of data read and written in strictly defined size increments; CosmosDB applies the same pricing to request units but does not explicitly define expected consumption; and Datastore has higher costs per operation but makes the cost independent of the item size. To understand the impact of price differences, we analyze the full execution of the Trip Booking benchmark. One workflow invocation requires three insertions and three deletions, with all items taking at most a few hundred bytes. While the estimated storage costs are similar on each platform, between ¢0.68 and ¢1.08 for one thousand executions, they impact the final cost differently. NoSQL operations add only 2.74% and 6.72% of the total price on AWS and GCP, respectively. The total execution cost on Azure is just ¢2.4. There, the estimated cost of CosmosDB request units is equal to ¢0.68 and adds 28.5% of workflow price.
@@
Figure 15: Price per 1000 workflow executions: function costs are opaque and state transition costs are translucent.
VII-F RQ5: Evolution of Performance
@@
Figure 16: Comparison of critical path (opaque) and overhead (hatched) between 2022 and 2024, burst invocations.
VII-G Threats to Validity
@@
A threat to the external validity is our choice of benchmark applications. We mitigate this by using applications from different domains that correspond to previous findings on the characterization of workflow use cases [41, 10]. Regarding internal validity, the different geographical regions and different week days we conducted our measurements on could have an impact. While we repeat each experiment six times to obtain stable results, there could be performance variability based on the time of day. However, systematically investigating this is beyond the scope of our work.
VIII Related Work
Multiple benchmark suites have been proposed to cover different aspects of serverless computing, from microarchitecture to the application level [59, 60, 11, 61, 62, 12, 63, 64, 65, 18, 66, 17, 67]. Kousiouris et al. [68] use microbenchmarks to estimate the overhead of orchestration in OpenWhisk. López et al. [69] investigate the orchestration overhead with microbenchmarks of function chains and parallel functions. Shahidi et al. [70] evaluate the performance and cost of two stateful workflows on AWS and Azure. Barcelona-Pons et al. [71] use a microbenchmark to test the performance of fork-join parallelism in workflow orchestrators. Other performance studies of serverless focus on non-workflow orchestration [72, 65, 73].

### v3

@@ 7.2. RQ1: Runtime Differences among Platforms
Figure 8. Runtime of benchmark applications on AWS Step Functions, GC Workflows, and Azure Durable, burst invocations.
7.3. RQ2: Causes for Runtime and Stability Differences
According to our results, AWS and Google Cloud provide a performance-reliable workflow service, whereas the variability is considerably higher on Azure. Thus, we split the runtime into two components to investigate the reasons behind this: the critical path $T_{C}$, computed as the sum of all states’ maximum runtime within one phase, and the overhead $T_{O}$ caused by the scheduling and data movement conducted by the cloud workflow service. We calculate the overhead $T_{O}$ by subtracting the critical path $T_{C}$ from the total runtime. Figure 9 shows the critical path and overhead for all benchmarks. Azure’s runtime is dominated by highly variable scheduling overhead: For example, the overhead of the ExCamera benchmark is, on average, 495.5s, more than 36$\times$ as long as its critical path of 13.5s. The ML benchmark incurs the least overhead of 5$\times$ the length of its critical path. Also, Azure’s critical path is very fast across all benchmarks, demonstrating the fastest critical path for ExCamera, MapReduce, and Machine Learning. Google Cloud, however, has the slowest critical path throughout the entire benchmark suite. In summary, orchestration overhead causes long runtimes and performance variances on Azure. For AWS and Google Cloud, however, the critical path varies. We therefore explore different causes for the differing behavior of the cloud platforms in the following Sections 7.3.1 and 7.3.2.
@@ 7.3. RQ2: Causes for Runtime and Stability Differences
Figure 9. Critical path (opaque) and overhead (hatched) of different benchmarks on considered platforms, burst invocations.
7.3.1. RQ2.1 Sources of Overhead
@@
The data downloaded from the object storage differs between benchmarks (Table 4, p. 4), with hundreds of megabytes in ExCamera, 1000Genomes, and Video Analysis. These benchmarks experience the highest relative overhead of 36.7$\times$, 10$\times$, and 14.95$\times$ their critical paths on Azure. To verify that this correlation is indeed causation, we execute a microbenchmark evaluating the cloud storage I/O performance. We invoke 20 functions in parallel where each attempts to download a file of size $D$ from the storage. Figure 10(a) shows that the overhead remains stagnant for AWS at around one second and nearly stagnant on Google Cloud, increasing a bit for downloads larger than 1MB. On Azure, however, we observe an overhead of almost 149 and 4.9 seconds for 128 and 1 MB files, respectively. Therefore, data downloads can account for a significant part of the large overhead measured on Azure Durable.
@@
(b) Invocation latency, chain of 10 functions, $2^{5}\leq M<2^{18}$, 256MB, warm invocations.
Figure 10. Analysis of different sources of overhead.
@@
Another potential source of overhead are parallel invocations within a benchmark: Benchmarks with the highest degree of parallelism – ExCamera and 1000Genomes – show the largest overheads of Azure. We test this by executing a microbenchmark that spawns $N$ functions in parallel, each one sleeping for $T$ seconds, and start 30 such invocations concurrently. Figure 11 shows the relative overhead of the actual runtime of the workflow compared to the function execution time. AWS functions demonstrate modest overhead, with largest values for the shortest duration. The absolute overhead incurred for a certain number of parallel functions remains relatively constant, causing the relative overhead to decrease with higher function executin times. GC functions present a larger relative slowdown that increases with the number of parallel tasks. There, the system puts a cap on scaling up and reuses containers, as 30 invocations with $N=2,T=1$ start 60 different function containers on AWS, but only 30 on Google Cloud. On the other hand, Azure experiences an order of magnitude larger relative overhead that increases with the parallelism factor but does not seem to be correlated to the function runtime.
@@
Figure 11. The overhead of parallel sleep microbenchmark, $2\leq N\leq 16$, $1\leq T\leq 20$, 256MB, burst invocations.
@@
Figure 12. Scaling profiles: the number of distinct containers used for 30 consecutive workflow invocations.
@@ Conclusions
The microbenchmarks demonstrate that a significant part of the overhead observed on Azure originates from the parallel schedules and storage I/O. Moreover, the return payload can be a source of overhead on Azure for larger payloads. To minimize overheads, workflows downloading large amounts of data, using high levels of parallelism, and high return payloads may therefore better be deployed to AWS or Google Cloud, with AWS demonstrating less overheads and better scalability across our benchmarks.
7.3.2. RQ2.2 Critical Path Discrepancy
@@
The cloud provider controls the CPU allocation to a serverless function, either in relation to the memory configuration on AWS and GCP (lambda-vcpu, 2024; Google Cloud, 2024a), or in an undisclosed fashion on Azure. We use the selfish detour benchmark to quantify OS noise (Hoefler et al., 2007), which allows us to estimate how long the function is suspended by the OS, which in turn approximates the vCPU timeshare. The benchmark runs a tight loop and records the event that one iteration took significantly more cycles than expected $N$ times. The magnitude and frequency of these events characterize the suspension and noise. We deploy a workflow with a single function executing the benchmark, invoke it 30 times concurrently, collect $N=5000$ events, and sample warm invocations to obtain consistent results. Figure 14(a) compares the relative to the expected suspension time according to the cloud documentation. We observe less noise on Google Cloud when compared to AWS, with more than 20% difference on 1024MB memory. We normalize the critical path per platform using the following approximation: given a function with memory configuration $M$, we represent the relative duration of function suspension as $S_{M}$ and compute the normalized critical path $T^{\prime}_{C}=T_{C}*\left(1-S_{M}\right)$. We observe the largest relative discrepancy on two benchmarks, MapReduce (Figure 14(b)) and Machine Learning (Figure 14(c)). The overall trend observed in Section 7.2 remains unchanged: Google Cloud demonstrates the longest critical path duration. The suspension time explains the shorter critical path on Azure as compared to AWS and GCP for benchmarks with low-memory configurations: Azure functions receive larger CPU allocations.
Cold Starts
@@
Cold invocations add significant overhead to the function execution (Copik et al., 2021). Table 5 shows the frequency of cold starts in our measurements, with cold starts identified using the containerID (see Section 4.3). Azure Durable performs significantly better, experiencing almost no cold starts, likely because function apps on Azure can hold many invocations concurrently (Copik et al., 2021). While the low scalability causes high orchestration overheads, it benefits the computations by putting them in warm containers. Figure 13 shows the impact of cold starts on the critical path and overhead. We show only the duration of warm starts on Azure Durable, as all benchmarks show a very low amount of cold starts on Azure. On AWS and GCP, however, there is a high percentage of cold starts in our measurement data. We therefore collected another 60 workflow invocations with at least one warm function and show the critical path for the resulting completely warm invocations. Google Cloud and AWS functions perform up to 2.0×2.0$\times$2.0 × and 4.5×4.5$\times$4.5 × better, respectively, achieving almost the same performance as Azure. Thus, cold starts are a major factor influencing the slowdown and performance instability observed in many benchmarks.
Conclusions
Our experiments show that Azure achieves short critical paths due to larger CPU allocations and less cold starts. To minimize the critical path, benchmarks using a low-memory configuration and only being executed occasionally can therefore be deployed to Azure. High-memory configurations, however, get higher CPU shares on AWS and GCP. GCP shows the slowest critical path even for warm invocations, while AWS can be competitive to Azure, making it a good choice for frequently executed workflows.
@@
Table 5. Relative #cold starts and #state transitions.
@@
Figure 13. Critical path (opaque) and overhead (hatched) of warm and cold invocations.
@@
Figure 14. Analysis of OS noise.
@@
Figure 15. Scalability of 1000Genome workflow.
7.4. RQ3: Usability for Scientific Workflows
@@
There is rising interest in the scientific community to use serverless solutions (Eismann et al., 2020), accompanied by experimentation with serverless offerings of the platforms (Malawski et al., 2020) and management systems for serverless execution of workflows (John et al., 2019; Jiang et al., 2017; Roy et al., 2022b, a). However, they do not consider the workflow orchestration systems the cloud platforms offer. We use the scientific benchmark 1000Genome to compare cloud services and the HPC system Ault using nodes equipped with Intel(R) 6154@3.00GHz CPU, repeating measurements five times.
@@
First, we compare the runtime of the total workflow, as shown in Figure 14(d). While the workflow execution time is, on average, 457.7s and 259.8s on GCP and AWS, respectively, the execution takes only 7.7s on Ault. GCP exhibits a coefficient of variation of 12.2%, while AWS has a coefficient of variation of only 3.3% - even lower than 4.1% on Ault. Interestingly, I/O takes less than one second on AWS, meaning that the computation is slower in the cloud. Then, we compare the scaling behavior of the different platforms for the individuals task of the workflow. We employ strong scaling, i.e., adding more jobs while keeping the size of the input file the same, resulting in smaller chunks per job. Figure 14(e) shows the speedup of 1.96 and 1.95 on AWS, 1.91 and 1.95 on GCP, and 1.51 and 1.24 on Ault for 10 and 20 jobs w.r.t. 5 and 10 jobs, respectively. The cloud platforms achieve a nearly-optimal speedup, which is not surprising given the high overhead for the baseline execution.
7.5. RQ4: Pricing
@@
We compare the average cost of executing a workflow and estimate the prices, as shown in Table 3, p. 3. Functions invoked during the execution of a workflow are billed based on the integral of memory and duration. Figure 16 visualizes the cost of workflow execution split into two groups: function execution (opaque) and the cost of orchestrating the state machine (hatched). Note that, due to Azure’s billing and measurement system, we could only retrieve an average cost value over all workflow invocations. Even though the Trip Booking benchmark is a simple pipeline with error catching, running it with workflow orchestration still adds significant state transition costs. Azure is the most expensive service for the 1000Genome benchmark. Google Cloud is the most expensive for MapReduce due to the high number of state transitions. AWS Step Functions are the most expensive solution for the other four benchmarks because functions cost 6.7×6.7$\times$6.7 × more for computation than Google Cloud Functions. The price charged for state transitions is nearly identical between AWS and Google Cloud, even though AWS charges 2.5×2.5$\times$2.5 × more: the AWS state language requires fewer states to implement the benchmarks (Table 5). Overall, we observe that Azure is expensive for workflows where it is also the slowest platform, such as for 1000Genome, but offers the cheapest pricing for benchmarks it also executes fastest, such as ML and MapReduce. Contrary to that, AWS shows high pricing for benchmarks it is fastest, such as Video Analysis and ExCamera
@@
In addition to execution and orchestration costs, workflows generate charges when accessing the object and NoSQL storage. While the prices of read and write operations on the object storage are the same across clouds, the billing models for key-value storage differ: DynamoDB charges according to the amount of data read and written in strictly defined size increments; CosmosDB applies the same pricing to request units but does not explicitly define expected consumption; and Datastore has higher costs per operation but makes the cost independent of the item size. To understand the impact of this, we analyze the full execution of the Trip Booking benchmark. One workflow invocation requires three insertions and three deletions, with all items taking at most a few hundred bytes. While the estimated storage costs are similar on each platform, between ¢0.68 and ¢1.08 per 1000 executions, they impact the final cost differently. NoSQL operations add only 2.74% and 6.72% of the total price on AWS and GCP, respectively. The total execution cost on Azure is just ¢2.4. There, the estimated cost of CosmosDB request units is equal to ¢0.68 and adds 28.5% of workflow price.
@@
Figure 16. Price per 1000 workflow executions: function costs are opaque and state transition costs are translucent.
7.6. RQ5: Evolution of Performance
@@
Figure 17. Comparison of critical path (opaque) and overhead (hatched) between 2022 and 2024, burst invocations.
7.7. Threats to Validity
@@
A threat to the external validity is our choice of benchmark applications. We mitigate this by using applications from different domains that correspond to previous findings on the characterization of workflow use cases (Eismann et al., 2020, 2022). Regarding internal validity, the different geographical regions and different week days we conducted our measurements on could have an impact. While we repeat each experiment six times to obtain stable results, there could be performance variability based on the time of day. However, systematically investigating this is beyond the scope of our work.
8. Related Work
Multiple benchmark suites have been proposed to cover different aspects of serverless computing, from microarchitecture to the application level (Ustiugov et al., 2021; Shahrad et al., 2019; Copik et al., 2021; faa, 2020; Back and Andrikopoulos, 2018; Kim and Lee, 2019; Maissen et al., 2020). However, all of them consider only the execution of single functions. Das et al. (2018) benchmark serverless edge computing platforms. Other performance studies of serverless applications focus on non-workflow orchestration systems, e.g., using cloud storage and queue triggers (Scheuner et al., 2022; Grambow et al., 2021; Hancock et al., 2022; Somu et al., 2020). Grambow et al. (2021) propose BeFaaS, providing an application benchmark modeling an online shop where the functions communicate using synchronous and asynchronous calls. In contrast, SeBS-Flow targets serverless workflow orchestrations.
ServerlessBench (Yu et al., 2020) considers a function chain microbenchmark orchestrated by AWS Step Functions, but only measures runtime of the workflow and time in between function invocations for varying payload sizes. Kousiouris et al. (Kousiouris et al., 2022) use microbenchmarks to estimate the overhead of orchestration in OpenWhisk. López et al. (García López et al., 2018) investigate the orchestration overhead with microbenchmarks of function chains and parallel functions. Shahidi et al. (Shahidi et al., 2021) evaluate the performance and cost of two stateful workflows on AWS and Azure. Barcelona-Pons et al. (Barcelona-Pons et al., 2019) use a microbenchmark to test the performance of fork-join parallelism in workflow orchestrators. With SeBS-Flow, we provide not only microbenchmarks, but also six applications from different domains that can automatically be deployed to different cloud platforms. Based on these benchmarks, we present a broader evaluation of the performance of cloud platforms.